### v1

# Adaptive Layer-skipping in Pre-trained LLMs

Abstract
Various layer-skipping methods have been proposed to accelerate token generation in large language models (LLMs). However, they have overlooked a fundamental question: How do computational demands vary across the generation of different tokens? In this work, we introduce FlexiDepth, a method that dynamically adjusts the number of Transformer layers used in text generation. By incorporating a plug-in router and adapter, FlexiDepth enables adaptive layer-skipping in LLMs without modifying their original parameters. Introducing FlexiDepth to Llama-3-8B model achieves layer skipping of 8 layers out of 32, and meanwhile maintains the full 100% benchmark performance. Experimental results with FlexiDepth demonstrate that computational demands in LLMs significantly vary based on token type. Specifically, generating repetitive tokens or fixed phrases requires fewer layers, whereas producing tokens involving computation or high uncertainty requires more layers. Interestingly, this adaptive allocation pattern aligns with human intuition. To advance research in this area, we open sourced FlexiDepth and a dataset documenting FlexiDepth’s layer allocation patterns for future exploration.
1 Introduction
Large language models (LLMs) have achieved remarkable success in various tasks, including translation [44], code generation [16], and math reasoning [6]. Currently, LLMs typically generate each token by performing a full forward pass through all Transformer decoder layers. However, such a uniform allocation is counter-intuitive, as simpler tasks intuitively require fewer computational resources, whereas more complex tasks demand more extensive processing. This uniform allocation not only results in computational inefficiencies but may also contribute to overfitting in LLMs. To improve the efficiency of token generation, various layer-skipping techniques have been proposed. One line of research leverages statistical information, such as the difference between layer inputs and outputs, to identify and skip less important layers [26, 13, 43]. Another collection of methods involves early-exit [37, 18, 34], where a confidence measure or a router dynamically determines whether to bypass all subsequent layers at an intermediate point of LLM. While these techniques have successfully reduced computational costs, they overlook a fundamental question: How do computational demands vary across the generation of different tokens
In this work, we propose FlexiDepth, a method for adaptive layer-skipping in pre-trained large language models. At each transformer layer, our method determines whether the hidden state input shall forward pass the layer or skip it. This layer-wise approach offers the flexibility to tailor the computation path for each token, enabling processing with varying numbers of transformer layers.
Contrast to layer skipping methods requiring training LLM from scratch, i.e. mixture-of-depth (MoD) [31], FlexiDepth enables adaptive layer-skipping in pre-trained LLMs without modifying their original parameters. Specifically, FlexiDepth introduces two light-weight plug-in modules at each decoder layer: (1) a router that makes binary decisions about whether to skip the layer, and (2) a adapter that resolves representation mis-alignments caused by layer skipping. The router and adapter are the only trainable components within our framework. All other parameters are inherited from a pre-trained LLM and remain frozen. At each layer, the input hidden states are first evaluated by the router and directed to go through or skip the layer. For hidden states that go through the layer, they are processed by the original attention and FFN modules of the LLM. Conversely, hidden states that skip the layer bypass these modules and are instead processed by the adapter. The adapter’s role is to transform these skipped hidden states, aligning their representation to match those undergo full processing. Finally, both the adapted hidden states and the processed hidden states are combined to produce the layer output.
Figure 1: Token depth map illustrating layer-skipping patterns when applying FlexiDepth to Llama-3-8B-Instruct. The light-to-dark blue gradient indicates layer usage ranging from 16 to 32.
To ensure compatibility with autoregressive generation, FlexiDepth computes the KV cache [28] for all hidden states, including those that skip this layer. This crucial step preserves the full context required for the attention mechanism. Without maintaining the KV cache, subsequent queries would be unable to attend to the hidden states that skipped the layer, rendering those tokens permanently unavailable for future generation. This would compromise the model’s ability to produce coherent and contextually accurate sequences. Preparing the full KV cache offers a simple method to maintain the integrity of autoregressive generation.
We comprehensively evaluate FlexiDepth across a diverse set of benchmarks. When applied to a Llama-3-8B-Instruct model [8], FlexiDepth preserves full performance (100.7%) while skipping 8 out of 32 layers, significantly outperforming existing layer-skipping methods, particularly in continuous generation tasks.
Beyond performance gain, FlexiDepth can reveal layer allocation patterns of pre-trained LLMs, addressing the question of how computational demands vary when generating different types of tokens. To demonstrate this phenomenon, we created a colored map that shows the number of layers used to generate each token, and call it "token depth map". This token depth map visualization, as shown in Figure 1 (left), reveals that for different language generation tasks, summarization typically utilizes more layers on average than extractive question answering. Similarly, in mathematical reasoning tasks like addition (Figure 1, right), tokens on the left-hand side of equations require fewer layers to generate compared to those on the right-hand side. This type of adaptive allocation appears to resonate with human intuition. For instance, in language tasks, humans typically exert more effort in summarization to condense the entire context into a concise form. In contrast, extractive question answering often relies on retrieval and extraction, making it easier to identify the answer. Likewise, in mathematical tasks, deriving the right-hand side of an equation involves computation that consumes more steps. To inspire further research along this direction, we compile a dataset documenting the layer allocation behaviors of FlexiDepth across various tasks.
Figure 2: The FlexiDepth layer. Left: Full-processing path where hidden states undergo the pre-trained attention and FFN modules. Right: Skipping path where hidden states bypass the attention module and processed by a lightweight adapter. The router and adaptor (in red) are the only trainable components within the FlexiDepth Block.
2 Method
In this section, we present the detailed implementation of FlexiDepth, which consists of multiple FlexiDepth layers. Each FlexiDepth layer enhances a standard pre-trained decoder layer with a plug-in router and adapter, allowing adaptive layer-skipping without modifying the original parameters. Different from Coda [21] and COLT5 [1] that adopt adapters for conditional computation in transformer encoders, our method focuses on seamless integration with decoder-only models. As illustrated in Figure 2, in each FlexiDepth layer, the router first processes normalized input hidden states to make skipping decisions. Based on the router’s output, each hidden state either follows the full-processing path (Figure 2, left) or the skipping path (Figure 2, right). The outputs from both paths are then combined to produce the layer output. In the following parts, we elaborate on the core designs of FlexiDepth, including router design, attention skipping, FFN skipping, and layer-skipping loss.
2.1 Router design
FlexiDepth employs a router to compute gating scores for routing. Let $X=[x_{1},x_{2},\ldots,x_{T}]\in\mathbb{R}^{T\times d}$ denote the input hidden states, where $T$ is the sequence length and $d$ is the hidden dimension. The router computes their corresponding gating scores $G=[g_{1},g_{2},\ldots,g_{T}]\in\mathbb{R}^{T}$ as follows:
$G=\sigma(\text{Router}(\text{Norm}(X))),$ (1)
where Norm refers to the RMSNorm [42], and $\sigma$ is the sigmoid function to ensure $g_{i}\in(0,1)$. In previous works [31, 36], the router is typically implemented as a simple linear transformation, and jointly optimized with the transformer during training. However, in our setting, the original parameters of the pre-trained transformer are frozen. Merely relying on a linear transformation is insufficient to capture the nuances of the hidden states for routing. To address this issue, we design a parameter-efficient router based on a bottlenecked MLP:
$\text{Router}(z)=W_{\text{r}}\cdot(W_{\uparrow}\cdot\text{Norm}(\tanh(W_{\downarrow}z))),$ (2)
where $W_{\downarrow}\in\mathbb{R}^{d_{r}\times d}$, $W_{\uparrow}\in\mathbb{R}^{d\times d_{r}}$, and $W_{\text{r}}\in\mathbb{R}^{1\times d}$ represent the weights of down-projection, up-projection, and router head, respectively. The $d_{r}$ denotes the bottleneck dimension.
Upon obtaining the gating scores $G$, we implement a threshold-based routing mechanism with a predefined threshold $\tau$. For each hidden state $x_{i}$, if $g_{i}>\tau$, it will be routed to the full-processing path (Figure 2, left), with the output multiplied by $g_{i}$ to maintain gradient flow to the router. Conversely, if $g_{i}\leq\tau$, it will be routed to the skipping path (Figure 2, right), with the output multiplied by $(1-g_{i})$.
2.2 Attention skipping
In the skipping path of a FlexiDepth layer (Figure 2 right), the input hidden states simply bypass the attention module via a shortcut. WhileHowever, this naive approach reduces computational overhead, it risks losing critical contextual informleads to a significant loss of contextual information, resulting in performance degradation. The attention mechanism depends onleverages query, key, and value vectors to enable tokens to attend to one another,each other, thereby capturing long-range dependencies essential for generation. When aFor hidden states that skips the attention module entirely, its, their corresponding key and value vectors are.. not generated generated, preventing subsequent tokens from accessing its informationttending to them. We show an example in Figure 3 to describe this.
As illustrated in Figure 3 (left), in the vanilla attention mask, every token can attend to all preceding tokens. For instance, query vector $q_{4}$ attends to keys $k_{0}$ through $k_{4}$. However, when employing FlexiDepth without maintaining the key-value cache (KV-cache), skipping hidden states such as $x_{1}$ and $x_{3}$ results in missing keys $k_{1}$ and $k_{3}$, as depicted in the attention mask labeled ’No KV Cache’ (Figure 3, middle). Consequently, future tokens like $q_{4}$ lose access to the contextual information stored in these hidden states.
To address this issue, we propose a simple solution: we still compute the KV-cache for hidden states that skip the attention module. Specifically, we still bypass the computation of query vectors (e.g., $q_{1}$ and $q_{3}$) and their associated scaled dot-product operations for the skipped states $x_{1}$ and $x_{3}$. However, we retain the calculation of their corresponding keys $k_{1},k_{3}$ and values $v_{1},v_{3}$, as demonstrated in the attention mask labeled ’KV Cache’ (Figure 3, right). This ensures that future generation maintains access to the entire context, preserving the integrity of autoregressive generation.
Figure 3: Comparison of attention masks. Left: the vanilla attention mask. Middle: the attention mask without KV Cache, where hidden states $x_{1}$ and $x_{3}$ skip this layer. Right: the attention mask with KV Cache, where despite $x_{1}$ and $x_{3}$ skipping this layer, their corresponding keys and values are still computed.
2.3 FFN skipping
We observe that directly skipping the feedforward network (FFN) via a simple shortcut leads to significantly degrade model performance. Unlike the attention module, which only consists of linear $q$, $k$, $v$ transformations, the FFN module introduces nonlinearities. Consequently, hidden states transformed by the FFN might not share the same latent space as those skipping it. To mitigate this issue, we employ a lightweight adapter to align their representations. This adapter follows the same structure as the FFN but features a significantly reduced intermediate dimension, with a reduction factor of 16x. As shown in Figure 2 right, the adapter is placed at the same position as the skipped FFN.
2.4 Layer-skipping loss
To balance computational efficiency and generation quality, we introduce a skipping loss that is jointly optimized with the next-token prediction loss. This loss is designed to penalize the squared sum of the layers used. The squared loss imposes a larger penalty on tokens that activate more layers and a smaller penalty on those that use fewer. This non-uniform penalty stabilizes training by preventing the model from falling into extreme patterns, such as skipping all layers or none. The loss function is formulated as:
$\mathcal{L}_{skip}=\frac{1}{T}\sum_{t=1}^{T}\left(\sum_{l=1}^{L}g^{l}_{t}\right)^{2},$ (3)
where $g^{l}_{t}$ denotes the gating score for layer $l$ at time step $t$. The final loss integrates the skipping loss $\mathcal{L}_{\text{skip}}$ with the original language modeling loss $\mathcal{L}_{\text{lm}}$, weighted by a factor $\alpha$:
$\mathcal{L}=\alpha\cdot\mathcal{L}_{\text{skip}}+\mathcal{L}_{\text{lm}}.$ (4)
3 Experiments
3.1 Implementation details
We implement FlexiDepth on the pre-trained Llama-3-8B-Instruct model [8], which consists of 32 transformer layers. To enable adaptive layer-skipping, we convert the latter 16 layers into FlexiDepth layers, as prior works [26, 35] and empirical studies suggest that skipping earlier layers degrades performance. For each FlexiDepth layer, the router, as defined in Equation 3, uses a bottleneck dimension $d_{r}=\frac{1}{16}d$, where $d$ is the hidden dimension. The router’s gating function is implemented using SparseMixer [23, 24] to ensure differentiability. The adapter has the same structure as the original FFN but reduces the intermediate dimension by a factor of 16. For the loss function in Equation 4, we set the coefficient $\alpha=1\times 10^{-3}$, which results in approximately 8 layers skipped during generation. We train FlexiDepth on the Tulu-v2 dataset [17] for 3 epochs using the AdamW [25] optimizer with a learning rate of $1\times 10^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=1\times 10^{-8}$. We use a warmup ratio of 0.03 and a global batch size of 64. The training takes approximately 7 hours on 8 NVIDIA A100-PCIE-40GB GPUs.
3.2 Main results
In this part, we present the experimental results of applying the proposed FlexiDepth to the pre-trained Llama-3-8B-Instruct model [8].
Benchmarks.
We evaluate FlexiDepth across a diverse set of tasks. For single-token generation benchmarks, we include: MMLU [14], HellaSwag [40], and Winogrande [33]. For multi-token generation tasks, we test on: GSM8K [5], HumanEval [4], CoQA [32]. Evaluations are conducted using the lm-evaluation-harness toolkit [10], with 5-shot settings for MMLU, HellaSwag, Winogrande, and GSM8K, and zero-shot settings for HumanEval and CoQA. The evaluation metrics are accuracy (acc) for MMLU, normalized accuracy (acc_norm) for HellaSwag, accuracy (acc) for Winogrande, exact-match (EM) for GSM8K, pass-at-1 (Pass@1) for HumanEval, and F1 score (F1) for CoQA.
Baselines.
We compared FlexiDepth with previous layer-skipping methods that are compatible with LLMs. All baseline methods are applied to the LLama-3-8B-Instruct model, which consists of 32 transformer layers. We define $k$ as the number of layers to skip during generation, with all methods configured to skip the same number of layers for fair comparison. The baseline methods are as follows: (1) LayerSkip [9]: This early-exit method skips the last $k$ consecutive layers during decoding and employs speculative decoding [22] to refine generation results. For comparison, we disable speculative decoding. (2) ShortGPT [26]: This approach prunes $k$ layers deemed less important by assessing their input-output difference. (3) LaCo [39]: This method employs a layer-collapse strategy to reduce the model by $k$ layers, merging subsequent layers into a prior one using a Reserving-Differences-while-Seeking-Common (RDSC) approach. Please read the original paper for the details. (4) MindSkip [12]: This method explores attention, FFN, and layer skipping via a simple linear router, finding that skipping FFN or entire layers leads to significant performance degradation. For comparison, we employ its layer skipping setting and train it on the same tulu-v2 [17] dataset for comparison.
Methods Single-Token Generation Multi-Token Generation Retain %
MMLU Hellaswag Winogrande GSM8K HumanEval CoQA
Vanilla 0.673 0.706 0.744 0.679 0.299 0.784 100.0%
Skip 4 Layers
LayerSkip 0.659 0.636 0.676 0.004 0.0 0.350 54.0%
ShortGPT 0.664 0.662 0.700 0.536 0.092 0.145 69.1%
LaCo 0.671 0.693 0.724 0.581 0.031 0.778 81.7%
MindSkip 0.664 0.698 0.722 0.378 0.189 0.720 84.2%
Ours 0.663 0.724 0.756 0.695 0.390 0.810 106.5%
Skip 8 Layers
LayerSkip 0.650 0.525 0.640 0.0 0.0 0.049 43.9%
ShortGPT 0.307 0.462 0.597 0.001 0.0 0.005 32.0%
LaCo 0.656 0.628 0.695 0.065 0.006 0.707 65.3%
MindSkip 0.602 0.650 0.646 0.039 0.024 0.620 60.2%
Ours 0.616 0.705 0.735 0.662 0.341 0.801 100.7%
Table 1: Performance comparison based on Llama-3-8B-Instruct, which consists of 32 layers. Retain % represents the percentage of average retained benchmark performance.
FlexiDepth demonstrates a varied number of skipped layers for different tasks. To compare fairly with baselines at a consistent number of skipped layers $k$, we trained multiple FlexiDepth models with varying penalty weights $\alpha$ and evaluated their performance. For each task, we report the results of the model that achieves the target number of skipped layers $k$. Later, we will highlight FlexiDepth’s adaptability by showing its performance across these tasks with a single, fixed $\alpha$.
Table 1 presents the comparison results. FlexiDepth consistently outperforms baseline methods, especially in multi-token generation tasks. For example, when skipping 8 layers, baseline methods retain reasonable performance on single-token generation benchmarks and the short-form question-answering task CoQA. However, they suffer from near-zero accuracy on GSM8K and HumanEval that requires longer reasoning. In contrast, FlexiDepth excels across both single-token and multi-token generation tasks, achieving an average performance retention of 100.7%.
Interestingly, FlexiDepth even slightly surpasses the original model’s performance on certain benchmarks, highlighting the potential benefits of adaptive depth. To further examine whether this improvement stems from training data or the mechanism itself, we compare FlexiDepth with the fully fine-tuned allenai/llama-3-tulu-2-8b model on huggingface, which is trained on the same dataset as FlexiDepth. Notably, this model achieves 0.554 on GSM8K and 0.354 on HumanEval, while FlexiDepth achieves 0.695 and 0.390 on the same tasks—even when skipping 4 layers. This result suggests that the performance gains are not solely attributable to the training data. We hypothesize that its adaptive layer-skipping mechanism may implicitly act as a form of regularization, by skipping less informative or noisy parameters during inference, thus enhancing generalization. Future work is needed to explore this possibility and investigate the broader potential of adaptive layer-skipping.
Model MMLU Hellaswag Winogrande GSM8K HumanEval CoQA Retain %
Llama-2-13B 0.492 0.727 0.714 0.236 0.102 0.790 100.0%
FlexiDepth 0.481 0.741 0.716 0.253 0.096 0.780 100.2%
[…]
Llama-3-8B 0.673 0.706 0.744 0.679 0.299 0.784 100.0%
FlexiDepth 0.663 0.743 0.756 0.657 0.323 0.803 102.1%
[…]
Qwen-2.5-3B 0.651 0.702 0.639 0.576 0.487 0.705 100.0%
FlexiDepth 0.643 0.701 0.679 0.583 0.476 0.741 101.5%
[…]
Table 2: Performance and skipping patterns of FlexiDepth on instruction-tuned models. Each model includes baseline performance (first row), FlexiDepth’s performance (second row), and average skipped layers (third row). The ’-Instruct’ are omitted for simplicity.
3.3 FlexiDepth on different language models
We evaluate FlexiDepth on language models of varying sizes. These models are trained under the identical configuration and dataset described in our implementation details. All of the models are instruction-tuned models. Larger models, specifically Llama-2-13B-Instruct and Llama-3-8B-Instruct, exhibit greater number of layers skipped compared to the smaller Qwen-2.5-3B-Instruct [38]. Concretely, the Llama-2-13B-Instruct and Llama-3-8B-Instruct skip approximately 7 and 6 layers respectively across various tasks, whereas the Qwen-2.5-3B-Instruct skips only about 1 to 2 layers. This pattern suggests that larger models inherently possess higher redundancy, thus allowing more aggressive layer skipping without significant performance deterioration. These observations highlight the substantial potential of adaptive layer skipping methods when applied to larger language models.
3.4 Layer allocation dataset
To investigate these layer-skipping patterns and their relationship to task complexity, we constructed a dataset using FlexiDepth with Llama-3-8B-Instruct as the base model. This dataset captures layer usage for token generation in two key domains: language comprehension and math reasoning. By recording the number of layers utilized per token, we aim to uncover how FlexiDepth dynamically adjusts its depth to meet varying task demands.
Text Generation.
For text generation tasks, we constructed a dataset by randomly sampling 100 paragraphs from the XSum test set [27], which is a collection of news articles. We evaluated FlexiDepth on three subtasks for each paragraph: copying, summarization, and continuation:
Please copy this paragraph: <paragraph>...</paragraph>
Directly output the copied paragraph here:
Please summarize this paragraph into a single sentence:
<paragraph>...</paragraph>
Directly output the summarized paragraph here:
Please continue writing this paragraph: <paragraph>...</paragraph>
Directly output the continued paragraph here:
We recorded the number of layers used per token across all outputs: copying averaged 21.95 layers (variance 6.05), summarization averaged 28.65 layers (variance 18.31), and continuation averaged 30.27 layers (variance 12.59). These results reveal a clear pattern: tasks requiring deeper contextual understanding, such as continuation and summarization, utilize more layers compared to the simpler copying task. The higher variance in summarization (18.31) and continuation (12.59) versus copying (6.05) suggests greater variability in layer allocation, likely reflecting the diverse cognitive demands of generating concise summaries or creative continuations. Detailed examples of these patterns are presented in Appendix A.
Math Reasoning.
For math reasoning, we created a dataset of 100 samples, each featuring a randomly generated list of 5–10 integers between 10 and 99. This controlled setup enables us to test basic arithmetic operations systematically. FlexiDepth was prompted to process each list with three subtasks: repeating the list, calculating the sum, and the product.
Please repeat the following list for 5 times: [...]
Please calculate the sum of numbers in the following list: [...]
Please calculate the product of numbers in the following list: [...]
The layer usage pattern across these three tasks was as follows: repeating averaged 20.09 layers (variance 8.08), addition averaged 22.45 layers (variance 14.67), and multiplication averaged 23.90 layers (variance 21.61). These findings align with task complexity—repetition, a straightforward memory task, uses the fewest layers, while multiplication, requiring iterative computation, demands the most. The higher variance in multiplication (21.61) compared to addition (14.67) and repetition (8.08) indicates that computational tokens exhibit more diverse layer needs, likely due to the increasing complexity of intermediate steps. Appendix A provides additional results, showing that tokens tied to arithmetic operations consistently require more layers than directly copying numbers from the context.
MMLU Hellaswag Winogrande GSM8K HumanEval CoQA Retain %
Vanilla 0.673 0.706 0.744 0.679 0.299 0.784 100.0%
FlexiDepth 0.663 0.743 0.756 0.657 0.323 0.803 102.1%
Linear Router 0.619 0.684 0.701 0.131 0.055 0.716 68.7%
No KV Cache 0.525 0.699 0.745 0.366 0.226 0.775 84.3%
No Adapter 0.226 0.356 0.579 0.004 0.0 0.047 28.1%
Table 3: Ablation studies based on Llama-3-8B-Instruct.
3.5 Ablation Studies
We show our ablation studies in this part. All experiments in this section utilize the same layer-skipping penalty, dataset, and training parameters as the original FlexiDepth implementation on the pre-trained Llama-3-8B-Instruct.
Linear Router.
In FlexiDepth, we employed a bottlenecked MLP as the router to facilitate nuanced routing decisions. When replacing this component with a simpler linear layer followed by a sigmoid activation, as previously used in mixture-of-depth (MoD) [31], we observe a substantial decrease in performance, particularly on the math reasoning task. Our analysis shows that the linear router mistakenly assigns too few layers to tokens requiring complex computations, significantly reducing the GSM8K performance from 0.657 to 0.131. Overall, the linear router retains only 78.6% of the original performance, emphasizing the importance of the bottlenecked MLP for routing.
No KV Cache.
We conducted an ablation study in which the KV cache is not computed for tokens routed to skip layers (illustrated as "No KV Cache" in Figure 3). Removing the KV cache results in substantial performance decline, retaining only 87.4% of the baseline performance. This highlights the need to retain contextual information during generation.
No adapter.
Many layer skipping methods do not have adapters. We therefore investigate the performance of FlexiDepth without adapter. This simplification drastically reduces model performance to merely 32.2% of the original FlexiDepth results. The adapter proves critical in aligning latent representations between skipped and processed states, ensuring representation coherence across the model’s layers.
4 Limitations
Although FlexiDepth reduces FLOPs by adaptively skipping transformer layers, our implementation does not lead to improved throughput on the existing GPU hardware. In FlexiDepth, samples within the same batch can take different execution paths during decoding. Consequently, each FlexiDepth layer must handle tokens that undergo full processing alongside those that skip layers, requiring simultaneous management of both computation and skip branches. This introduces overhead from control-flow management and irregular memory accesses, outweighing the speedup of theoretical FLOP reductions. Future work could investigate hardware-oriented optimizations to address these bottlenecks, leveraging techniques such as token grouping [30], expert sharding [30], and load balancing [15] to better align with GPU architectures. Such hardware-tailored enhancements could further unlock the efficiency potential of FlexiDepth in practice.
5 Related works
The existing language models can be broadly classified into three categories: encoder-only models (e.g., BERT [7]), encoder-decoder models (e.g., T5 [29]), and decoder-only models (e.g., GPT [3]). Across these architectures, various conditional computation methods have been proposed to optimize resource allocation by dynamically processing tokens.
In encoder-only models, conditional token processing has been explored to prune less critical tokens during processing. For instance, PoWER-BERT [11] and LTP [19] utilize attention scores to assess token importance, dynamically eliminating those deemed less significant. Similarly, LoT [20] introduces a router at each layer to determine whether tokens should be processed or skipped, tailoring computation to token relevance. These approaches leverage the bidirectional nature of encoders, allowing noncausal routing decisions based on full sequence context.
For encoder-decoder models, methods such as CoDA [21] and COLT5 [1] enable conditional computation within the encoder component. By replacing standard attention or feed-forward network (FFN) modules with lightweight alternatives for certain tokens, these techniques assign varying computational resources based on token complexity. However, their noncausal design constrains their applicability to decoder.
In decoder-only models, the autoregressive nature poses unique challenges for dynamic token processing, as tokens must be generated sequentially while preserving contextual dependencies. Piorneering efforts like MoD [31] and SkipLayer [41] address this by deploying a router at each Transformer layer, enabling tokens to dynamically bypass it. These models are trained end-to-end from scratch, integrating the routing mechanism into the learning process. Duo-LLM [2] also employs a router for conditional computation, but with a distinct method. It first constructs a dataset of oracle routing strategies via exhaustive enumeration and then trains the router on this data. While effective, these methods demand significant computational resources for training and dataset preparation, making them impractical for adapting pre-trained models. MindSkip [12] enables skipping in pre-trained models via learnable routers. It finds that bypassing FFN modules or entire layers significantly degrades performance and instead proposes a method to skip the attention module efficiently. Unlike these approaches, our method, FlexiDepth, enables adaptive layer-skipping in pre-trained models without compromising generation performance by using a plug-in router and adapter architecture.
6 Conclusion
We present FlexiDepth, an approach to dynamically skip layers in pre-trained large language models. It achieves state-of-the-art performance without altering the original model parameters. FlexiDepth offers valuable insights into the varying computational demands of token generation. We constructed a dataset capturing these allocation patterns to encourage further study of computational needs of tokens, thereby advancing the development of more adaptive and resource-efficient language models.
Appendix A Layer allocation Examples
In this section, we present examples to illustrate the layer allocation patterns of FlexiDepth, using the pre-trained Llama-3-8B-Instruct model [8] as the base.
A.1 Language Task
Copy:
Here is an example where FlexiDepth uses 21.84 layers in average per token:
Please copy this paragraph: $<$paragraph$>$Bissouma penned a professional deal with the northern club only in July 2016, four months after joining the reserves from AS Real Bamako. The 20-year-old, whose previous deal ran until June 2019, says he is very happy at Stade Pierre-Mauroy. "In football, when you put in the work you always get rewarded," Bissouma told the club’s website. "I still have a lot of progress to make. I want to give myself every chance of doing that and fulfilling my objectives. "The first is to keep the club in Ligue 1, and then I would also like to win a trophy with Lille." Bissouma has scored once in 17 appearances for the club since making his debut against Toulouse in September 2016.$<$/paragraph$>$ Directly output the copied paragraph here:
[…]
This task requires minimal contextual understanding, as it involves direct reproduction, which explains the relatively low layer usage. The first token ’B’ requires much more layers to generate because the generation of the first token involves the prefilling stage, where the whole context is been processed.
Summarization:
Here is an example using 29.15 layers in average per token:
Please summarize this paragraph into a single sentence: $<$paragraph$>$Bissouma penned a professional deal with the northern club only in July 2016, four months after joining the reserves from AS Real Bamako. The 20-year-old, whose previous deal ran until June 2019, says he is very happy at Stade Pierre-Mauroy. "In football, when you put in the work you always get rewarded," Bissouma told the club’s website. "I still have a lot of progress to make. I want to give myself every chance of doing that and fulfilling my objectives. "The first is to keep the club in Ligue 1, and then I would also like to win a trophy with Lille." Bissouma has scored once in 17 appearances for the club since making his debut against Toulouse in September 2016.$<$/paragraph$>$ Directly output the summarized paragraph here:
Bissouma, who signed a professional deal with Lille in July2016 and has scored once in17 appearances, is happy at the club and aims to keep them in Ligue1 and win a trophy.<|eot_id|>
Summarization involves understanding and condensing the input into a concise form, requiring deeper contextual abstraction and rephrasing, which leads to a higher layer usage of 29.15 compared to copying. We observe that tokens forming predictable collocations, such as "souma" following "Bis" in "Bissouma" or "trophy" after "win a" in "win a trophy," require fewer layers due to their low uncertainty after the initial tokens are generated.
Continuation:
Here is an example that utilizes 31.14 layers in average:
Please continue writing this paragraph: $<$paragraph$>$Bissouma penned a professional deal with the northern club only in July 2016, four months after joining the reserves from AS Real Bamako. The 20-year-old, whose previous deal ran until June 2019, says he is very happy at Stade Pierre-Mauroy. "In football, when you put in the work you always get rewarded," Bissouma told the club’s website. "I still have a lot of progress to make. I want to give myself every chance of doing that and fulfilling my objectives. "The first is to keep the club in Ligue 1, and then I would also like to win a trophy with Lille." Bissouma has scored once in 17 appearances for the club since making his debut against Toulouse in September 2016.$<$/paragraph$>$ Directly output the continued paragraph here:
Despite his young age, Bissouma has already shown impressive maturity in his performances, with his manager, Marcel Blondeau, praising him for his ability to adapt to the teams system. "He is a very intelligent player," Blondeau said. "He can play in different positions and has a good understanding of the game. He is also very strong in the tackle and can win the ball back." With his contract running until2021, Bissouma is expected to continue playing a crucial role in Lilles midfield for the coming seasons.<|eot_id|>
Continuation demands creative generation and the maintenance of contextual coherence, necessitating long-range dependencies and novel content creation, which results in a higher layer usage of 31.14. The generation process typically engages more layers, except for predictable tokens like names ("Bissouma" and "Blondeau"), which require fewer layers due to their memorized nature and low uncertainty.
A.2 Math Task
Repeating:
Here is an example that uses 20.15 layers in average:
Please repeat the following list for 5 times: [63, 15, 43, 75, 72, 61, 48, 71]
Here is the list repeated5 times
[63,15,43,75,72,61,48,71]
[63,15,43,75,72,61,48,71]
[63,15,43,75,72,61,48,71]
[63,15,43,75,72,61,48,71]
[63,15,43,75,72,61,48,71]<|eot_id|>
Repeating is a memory-based task requiring minimal computation, as it involves recalling and reproducing the input list, which explains the low layer usage. This example fits within the reported variance (8.08), reflecting the task’s consistent layer needs.
Addition:
Here is an example that uses 22.31 layers in average:
Please calculate the sum of numbers in the following list: [63, 15, 43, 75, 72, 61, 48, 71]
Lets calculate the sum of the numbers in the list
63 +15 =78
78 +43 =121
121 +75 =196
196 +72 =268
268 +61 =329
329 +48 =377
377 +71 =448
The sum of the numbers in the list is 448<|eot_id|>
Addition requires iterative computation, using more layers than repeating since FlexiDepth calculates each intermediate sum. The right-hand side of each equation (e.g., "78", "121") demands more layers due to the transformer’s internal computation, while the left-hand side (e.g., "63 + 15") generally uses fewer layers as it copies numbers from the input. We also find that the initial tokens on the left-hand side, such as "63" in "63 + 15", use slightly more layers than the rest of the left-hand side tokens. We hypothesize that this is to determine which intermediate result to copy from the context.
Production:
Here is an example that uses 24.51 layers in average:
Please calculate the product of numbers in the following list: [63, 15, 43, 75, 72, 61, 48, 71] To calculate the product of the numbers in the given list, Ill multiply each number individually
Heres the step-by-step calculation
1.63 x15 =945
2.945 x43 =40515
3.40515 x75 =3026125
4.3026125 x72 =218171500
5.218171500 x61 =13311439500
6.13311439500 x48 =637385920000
7.637385920000 x71 =45213914128000
The final product of the numbers in the provided list is:45213914128000<|eot_id|>
Multiplication involves complex iterative computations, as FlexiDepth handles large intermediate results and maintains numerical precision, resulting in higher layer usage. The right-hand side of each equation (e.g., "945", "40515") consistently uses the full number of layers to compute the product, reflecting the task’s high computational demand. Similar to addition, on the left-hand side, initial tokens like "63 × 15" in the first step may use slightly more layers to identify the intermediate result to copy from the context. Subsequent tokens (e.g., "945 × 43", "40515 × 75") exhibit a gradually decreasing trend in layer usage, possibly due to increasing certainty when retrieving intermediate results as the computation progresses.